If the line h passes through the two points (1, 2) and (3, 6) What is the slope of the line h passing through the two points?
Answer by dkppathak(439)

slop = m =y2-y1/x2-x1
given (x1,y1)= (1, 2)
(x2,y2)= (3, 6)
m = 6-2/3-1
m=4/2
m=2
slope can be written in the form of Tan x also